OpenStudy (anonymous):

Find the slope of a line that has an x-intercept of -2 and a y-intercept of 6.

OpenStudy (anonymous):

can anyone help me?

OpenStudy (anonymous):

anyone.......please i dont understand this question :(

OpenStudy (anonymous):

hi....

OpenStudy (anonymous):

hi...finally

OpenStudy (anonymous):

:) so for finding a slope we need 2 points

OpenStudy (anonymous):

do you knoiw these points?

OpenStudy (anonymous):

no, it didnt give me a graph

OpenStudy (ahsome):

Is there any other info we could sue?

OpenStudy (anonymous):

the line intercept with x at -2 which means ?

OpenStudy (anonymous):

at x=-2 y =0; so this is your first point

OpenStudy (anonymous):

second point is y=6 when x=0

OpenStudy (anonymous):

oh..so it would be (-2,0) and (6,0)?

OpenStudy (anonymous):

sorry i meant (0,6)

OpenStudy (anonymous):

\[m=\frac{ (y1-y2) }{ (x1-x2) }\]

OpenStudy (anonymous):

yes

OpenStudy (ahsome):

The standard Equation: \(y=mx+b\) We know that the \(y-intercept\) is equal to \(b\), so: \(y=mx+6\) We also know one point, the x-intercept. This is equal to \(-2,0\). Sub that into the equation: \(0=m*-2+6\) \(0=-2m+6\) Solve for the equation: \(0=-2m+6\) \(-6=-2m\) \(m=\frac{-6}{-2}\) \(m=3\)

OpenStudy (anonymous):

so which formula am i supposed to use....AJ01 or ahsome?

OpenStudy (ahsome):

Which ever makes sense for you @Addicto_

OpenStudy (anonymous):

so their both right?

OpenStudy (ahsome):

The equation: \[m=\frac{y_1-y_2}{x_1-x_2}\] Is more versatile. What I did worked for this question

OpenStudy (ahsome):

And yes, they both give the same answer

OpenStudy (anonymous):

ok thanks alot

OpenStudy (ahsome):

No problem @Addicto_ :)